# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v1

@@ 4. Future Trends
In this study, we summrized the relevant research on Android performance analysis and illuminated real-world challenges in Sections 2 and 3. Drawing upon these findings, we now present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
@@ 4.1. Targeted Objective Imbalance
Through our literature review (Section 2) and real-world exploration (Section 3), we identified a total of six types of performance issues and 40 factors, as shown in Fig. 5, where colors and white distinguish whether there is a relationship between factors and performance issues. We observed that, apart from the performance issue “Data Consumption”, all the other five performance issues involve corresponding factors, suggesting data consumption is an important but untouched topic.
In addition, of the 40 discovered factors, 21 have been analyzed in the literature, while the remaining 19 (almost half) have not been examined by our researchers, highlighting a significant research gap. Moreover, it’s intriguing to note that there are 5 factors (namely, Multi-Threads, Obsolete Task, Redundant Frames, SQL Statement, and Wakelock) that solely appear in the literature without a corresponding issue or complaint in real-world scenarios. This suggests that these factors may stem more from the assumptions of academic researchers rather than being genuine concerns of users or developers in practical settings.
The presence of this imbalance underscores the need for further research into issues genuinely concerning users or developers. Indeed, more effort should be put into addressing the untargeted real-world factors, aiming to bridge the existing gap between identified factors and practical concerns.
@@ 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our review also revealed that only 7 out of 40 factors (17.50%) are covered by state-of-the-art (SOTA) tools. We utilize Fig. 5 to further highlight this significant research gap, where Activity is the most targeted factor, followed by Resource and AsyncTask, and finally Service, Database Operation, Image, and Loop. Also, the majority of these tools are focused on issue detection, while only 3 of them (30.00%) are dedicated to fixing such issues. Although static, dynamic, and DL techniques have been demonstrated to be effective at analyzing performance issues in a series of research experiments, applying these approaches to reveal performance issues holistically in practice remains unsolved. Future research should not only focus on proposing a comprehensive framework to tackle many realistic challenges, such as API misuse, redundant objects, etc, which may require knowledge of code patterns as presented in Section 3.2.2, but also on finding solutions for fixing problems.
In addition, the adopted techniques are also highly imbalanced (most existing tools use traditional static or dynamic analysis, with only one work (Jabbarvand2020AutomatedCO, ) leveraging DL methods). The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e. code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Lu2021CodeXGLUEAM, ; Wang2022TestDrivenML, )). Thus, it is promising to apply advanced DL techniques to assist in solving realistic and specific performance issues. For example, convert code exhibiting the Unreleased References pattern mentioned in Section 3.2.2 into program analysis graphs (e.g., CFG and DFG). Use these graphs as features to train or finetune a GNN-based model (Zhou2018GraphNN, ; Xu2018HowPA, ; Velickovic2017GraphAN, ) to identify this anti-pattern. Leverage the network’s interpretability (Ying2019GNNExplainerGE, ; Yuan2021OnEO, ) to pinpoint problematic code segments and suggest optimizations, such as adding code to release references.
@@ 4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 11 (27.50%) of the 40 factors are covered by existing datasets. Even among the 11 factors, the data distribution is highly imbalanced. The number of activity samples is much larger than that of the other factors, which are more severe from a real-world perspective. In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Moreover, another major issue is the limited reproducibility. Our review indicates that only a small number of primary studies share their datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples.
Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible. The code examples we summarized from the exploratory study can all be used to construct a comprehensive dataset, covering 35 factors. Another possible solution involves leveraging the findings of our exploratory study, wherein we identify six specific code patterns from real-world scenarios. These patterns can act as seeds for mutation tools, enabling the generation of additional data to construct datasets.
@@ 5. THREATS TO VALIDITY
Our study has several factors that could affect the validity of our findings in terms of literature review and large-scale exploration.
For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 2.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant works, we conducted backward and forward snowballing and cross-checked our results.
For real-world data collection, the data filtering process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data.
@@ 6. Conclusion
In this paper, we conducted a comprehensive empirical study to reveal the current state of Android performance analysis and the research gap between real-world challenges and existing techniques. Based on our in-depth analysis of 48 relevant studies and real-world issues across three major platforms (Google Play, Stack Overflow, GitHub), we discovered that (1) There are 6 types of performance issues and corresponding 40 factors; (2) Performance issues are sophisticated, with a single issue potentially stemming from multiple factors and a single factor causing multiple issues; (3) The root causes of data consumption issue are not clear; (4) The primary concerns of researchers, developers, and users are not aligned; (5) The majority of existing approaches primarily utilize static analysis; (6) There is a significant gap in existing tools’ ability to address all factors causing performance issues; (7) There is an urgent need to construct benchmark dataset to evaluate automatic tools; (8) DL-based and LLM-based approaches have shown promising results in addressing software engineering problems, suggesting their potential for tackling Android performance issues as well.

### v2

@@ 4. Future Trends
Through our real-world exploration (Section 2) and literature review (Section 3), we have revealed significant gaps in current academic research: it addresses only a fraction of the performance issues that developers and users prioritize, existing tools lack the capacity to comprehensively tackle these issues, and current datasets are insufficient for effectively evaluating tool performance. Drawing upon these findings, we present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
@@ 4.1. Targeted Objective Imbalance
In Section 3.2.1, we found that there is a misalignment in the primary concerns of researchers, developers, and users: researchers primarily focus on energy consumption, developers prioritize memory consumption, while users are most concerned with responsiveness. This suggests that future research should increase focus on responsiveness and memory issues, as these have a more immediate impact on both end-users and developers. We also found two categories of performance issues—storage consumption and internet data usage—that remain unexplored in academic research, indicating potential areas for further investigation. Additionally, only 42.86% of the factors contributing to performance issues observed in real-world settings have been explored in academic research, remaining nearly 60% still awaiting for researchers to investigate. Interestingly, there are 23 factors among the 72 factors lead to performance issues are only been explored in literature but not observed in the real-world scenarios. This suggests that the issues currently discussed in academic research may not be of real concern to users or developers in real-world settings. Overall, this imbalance between academic research objectives and real-world challenges underscores the need for greater focus on the real-world performance issues and contributing factors that are most relevant to end-users and developers.
@@ 4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our study revealed that only 17 out of 72 factors (23.61%) are addressed by state-of-the-art (SOTA) tools, with most tools focusing on detection rather than resolution. In fact, only two out of the existing tools are dedicated to fixing performance issues. Although static, dynamic, and deep learning (DL) techniques have shown effectiveness in research experiments, applying these methods to holistically detect and fix performance issues in practice remains an unsolved challenge. Future research should focus not only on building comprehensive frameworks to detect real-world performance issues, but also on providing solutions for fixing these issues.
The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e. code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Wang2022TestDrivenML, )). Applying advanced DL techniques to detect and fix Android performance issues is promising, although significant challenges remain in collecting sufficient performance issue dataset for model training. The emergence of large language models (LLMs) also presents new possibilities. LLMs, with their in-context learning capabilities, could address the limitations of traditional DL approaches by enabling models to detect and resolve performance issues with fewer examples and greater adaptability (Hou2023LargeLM, ; Liao2023A3CodGenAR, ; Si2024AST, ; Fan2023LargeLM, ). This opens an exciting direction for future research in Android performance analysis, promising to bridge the current gaps in existing techniques.
@@ 4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 24 (33.33%) of the 72 factors are covered by existing datasets. Even among the 24 factors, the data distribution is highly imbalanced, as the smallest dataset contains only five apps, while the largest has 324. In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Another critical issue is reproducibility. Our study indicates that only a small number of primary studies share valid datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples. Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible.
@@ 5. THREATS TO VALIDITY
Our study has several factors that could affect the validity of our findings in terms of large-scale exploration and literature review. For real-world data collection, the data filtering process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data.
For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 3.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant work, we conducted backward and forward snowballing and cross-checked our results.
@@ 6. Conclusion
In this paper, we conducted a comparative study to examine real-world Android performance issues encountered by users and developers, and to evaluate the extent to which current academic research addresses these practical challenges. By analyzing large-scale discussions from three popular crowd-sourced platforms on Android performance issues, alongside a review of 66 high-quality research papers, we identified several gaps between real-world challenges and the focus of academic research. Our main findings are as follows: (1) A taxonomy categorizes seven types of performance issues and 72 contributing factors. (2) Six common code patterns of these performance issues. (3) There is a misalignment in the primary concerns of researchers, developers, and users. (4) Only 42.86% of the real-world factors have been explored in academic research. (5) The majority of existing tools primarily utilize static analysis, dynamic analysis, and deep learning, with most tools detecting issues rather than fixing them. (6) Existing available tools target only 23.61% of identified factors. (7) Existing available datasets cover only 33.33% factors and lack sufficient data to support effective tool evaluation. Finally, we proposed future research directions to bridge the identified gaps.